OpenStudy (anonymous):

factor 2x^2 + x -3 and 2x^2 + 5x + 3

OpenStudy (diyadiya):

have you tried?

OpenStudy (anonymous):

i just really dont know like whent theres a number in the front of the x

OpenStudy (diyadiya):

Ok, You're factoring by splitting the middle terms right? or Quadratic formula?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think splitting the middle terms

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i think this is calculus II

OpenStudy (anonymous):

any quadratic eqn is in the forum of ax^2 + bx + c = 0, in order to split the middle term ie bx just take the coeficient b and a and the constant term c, then think of any two numbers that when you add you get b and when you multply you get ac. in our first case 2x^2 + x - 3 =0, a =2, b = 1 and c = -3. therefore ac = -6 and b = 1 now numbers to be used are 3 and -2 ie the qn bcomes 2x^2 +3x - 2x -3 = x(2x+3)-1(2x+3) = (2x+3)(x-1)

OpenStudy (anonymous):

the second case 2x^2 + 5x + 3, here ac = 2x3 = 6 and b 5 therefore the numbers to be used to split 5x are 3 and 2 since 2+3 = 5 and 2x3 =6 therefore the question becomes 2x^2 +2x+3x+3 = 2x(x+1)+ 3(x+1) = (2x+3)(x+1)